# From Fuzzy to Exact: The Halo Architecture for Infinite-Depth Reasoning via Rational Arithmetic

### v2

@@ 6.4 Beyond Language: Implications for Time-Series Forecasting
While this work focuses on AGI reasoning, the "Exactness Hypothesis" generalizes to all recursive dynamical systems, particularly **Time-Series Forecasting**.
@@
6.5 Toward Verified AGI: A Formal Verification Roadmap
@@ Acknowledgements.
Special thanks to the open-source community for our simulation framework.
@@
This appendix provides the formal proofs for the theoretical bounds presented in the main paper, along with detailed hyperparameters to ensure reproducibility of the "Golden Triangle" experiments. The accompanying code submission (halo_proof_of_concept.py) contains the functional implementation of the Exact Inference Unit (EIU).
Appendix A Formal Proof of Theorem 1
Theorem 1 (The Halo Boundedness Theorem). Let $B_{ring}$ be the fixed bit-width of the Ring’s codebook, and $\alpha$ be the maximum bit-growth rate per layer (for standard matrix multiplication, $\alpha\approx 2$). Let $K$ be the Ring Reset Interval. For any infinite inference depth $L\to\infty$, the maximum bit-width $B_{max}$ required by the EIU is strictly bounded by:
@@ Appendix A Formal Proof of Theorem 1
Let $b_{t}$ denote the bit-width of the rational state representation $H_{t}$ at time step $t$. We model the inference process as a hybrid dynamical system with two modes of operation: Accumulation (The Light) and Collapse (The Ring).
1. Accumulation Phase (The Light): For any step $t$ where $t\not\equiv 0\pmod{K}$, the system operates in exact rational arithmetic. Given two rational numbers $q_{1}=n_{1}/d_{1}$ and $q_{2}=n_{2}/d_{2}$ with bit-widths $b_{1}$ and $b_{2}$, their product yields a new denominator $d^{\prime}=d_{1}d_{2}$. The bit-width of the product grows additively:
$\text{bits}(q_{1}\cdot q_{2})\approx\text{bits}(q_{1})+\text{bits}(q_{2})$
In a linear layer, this growth is bounded by the factor $\alpha$. Thus, the recurrence relation for bit-growth during the accumulation phase is:
$b_{t}=b_{t-1}+\alpha\quad\text{for }t\not\equiv 0\pmod{K}$ (16)
2. Collapse Phase (The Ring): At every $K$-th step ($t=mK$), the Ring Mechanism projects the continuous state $H_{t}$ onto the nearest valid rational in the fixed codebook $\mathbb{Q}_{grid}$. By definition, any element $q\in\mathbb{Q}_{grid}$ has a maximum bit-width of $B_{ring}$. Thus:
$b_{t}=B_{ring}\quad\text{for }t\equiv 0\pmod{K}$ (17)
3. Bounding the Maximum: Consider the interval between two resets, $t\in[mK,(m+1)K]$. The bit-width initializes at $B_{ring}$ at step $mK$. It then grows linearly for K1K-1 steps. The maximum width occurs at the step immediately preceding the next reset, i.e., $t_{peak}=(m+1)K-1$. Substituting into the recurrence:
$b_{peak}$ $=b_{mK}+\sum_{i=1}^{K-1}\alpha$
$b_{peak}$ $=B_{ring}+(K-1)\alpha$
Since the reset operation at $(m+1)K$ forces the width back to $B_{ring}$, $b_{t}$ never exceeds $b_{peak}$. Thus, the memory complexity is bounded by $\mathcal{O}(K)$ and is independent of the total depth $L$. ∎
@@ D.3 Silicon Feasibility: Complexity Reduction
Appendix E Appendix D: The Computational Tractability of Exactness
E.1 Architectural Mismatch of Commodity Hardware
  1. 1.
  2. 2.
  3. 3.
E.2 Optimizing Throughput: The EIU Solution
1. Lazy Reduction Strategy
2. "Fire-and-Forget" Asynchronous Pipelining
3. Hardware Binary GCD (Stein’s Algorithm)
4. Latency Analysis of Non-Linear Operations
Rational arithmetic handles non-linearities efficiently:
E.3 Silicon Feasibility: Complexity Reduction
Appendix F Appendix E: Training Dynamics and Differentiability
@@
F.1 Strategy 1: Handling Discreteness (VQ-VAE Paradigm)
@@ 2. Straight-Through Estimator (STE)
F.2 Strategy 2: Rational Mixed-Precision Training
@@
F.3 Transfer Learning: Post-Training Rationalization

### v3

@@ 6.3 Beyond Language: Implications for Time-Series Forecasting
While this work focuses on AGI reasoning, the "Exactness Hypothesis" generalizes to all recursive dynamical systems, particularly Time-Series Forecasting.
@@
6.4 The Ontological Necessity of Discretization
A profound theoretical objection to the Halo Architecture is the ubiquity of continuous representations (manifolds) in modern Deep Learning. Critics might argue that restricting AGI to the rational field $\mathbb{Q}$ fundamentally limits its expressivity. We challenge this view, positing instead that Discretization is the fundamental premise of Intelligence.
1. Semantic Crystallization via Phase Transition
In the continuous domain ($\mathbb{R}$), neural states are fluid and amorphous, susceptible to mixing and dilution. The Ring mechanism serves as a vehicle for Semantic Crystallization. It imposes a Phase Transition: forcing the fluid, probabilistic state to “snap” into a rigid, lattice-like logical structure. Without this periodic collapse, the reasoning chain evaporates into a superposition of meaningless values, a phenomenon we term Semantic Decoherence. Just as biological neurons fire in discrete spikes (All-or-Nothing) to transmit signals over long axons without attenuation, Halo uses discrete rationals to transmit logic over infinite depth.
2. The Quantum Error Correction Analogy
This process is structurally isomorphic to Quantum Error Correction. In quantum computing, a qubit maintains a superposition state but accumulates environmental noise (decoherence). To preserve information, the system must periodically measure the qubit (collapse the wavefunction) to correct errors before they become unrecoverable. Similarly, a "Thought" in an AGI system must be discretized to survive. The Ring acts as this measurement operator, collapsing the noisy continuous vector onto the nearest valid rational concept, effectively resetting the system’s entropy to zero.
3. The "Continuum Illusion"
Finally, we address the concern regarding transcendental numbers (e.g., $\pi$). Standard intuition assumes the physical world is a continuous differentiable manifold. However, fundamental physics (Bekenstein bound, Planck length) suggests that the universe is ultimately discrete and finite. The "smoothness" of real numbers is likely a macroscopic approximation of discrete states. By building AGI on $\mathbb{Q}$, we are not degrading reality; we are aligning our computational substrate with the discrete nature of information itself. Halo suggests that the "Unreasonable Effectiveness of Mathematics" arises not from continuity, but from the rigorous discrete structures (groups, rings, fields) that underpin it.
6.5 Toward Verified Infrastructure: A Formal Verification Roadmap
@@ Acknowledgements.
We thank Qiao Gao and Hao Wu for their constructive review and valuable feedback on the manuscript.
@@
Appendix A Appendix A: Formal Proofs of Theoretical Bounds
This appendix provides the rigorous mathematical derivations for the theoretical bounds presented in the main paper. We address two critical dimensions of the Halo Architecture: the boundedness of memory consumption (Theorem 1) and the superior numerical precision of the Ring mechanism (Theorem 2).
A.1  Formal Proof of Theorem 1 (The Halo Boundedness Theorem)
We first prove that despite the linearly growing precision of rational arithmetic in the "Light" phase, the "Ring" mechanism ensures the system’s memory complexity remains constant relative to depth.
Theorem 1 (Memory Boundedness). Let $B_{ring}$ be the fixed bit-width of the Ring’s codebook, and $\alpha$ be the maximum bit-growth rate per layer. Let $K$ be the Ring Reset Interval. For any infinite inference depth $L\rightarrow\infty$, the maximum bit-width $B_{max}$ required by the EIU is strictly bounded by:
@@ A.1  Formal Proof of Theorem 1 (The Halo Boundedness Theorem)
Let $b_{t}$ denote the bit-width of the rational state representation $H_{t}$ at time step $t$. We model the inference process as a hybrid dynamical system with two modes:
  1. 1.
    Accumulation Phase (The Light): For any step $t\not\equiv 0\pmod{K}$, the system operates in exact rational arithmetic. The bit-width grows additively based on the operation overhead $\alpha$:
    $b_{t}=b_{t-1}+\alpha$
  2. 2.
    Collapse Phase (The Ring): At every $K$-th step ($t=mK$), the Ring Mechanism projects the state onto the fixed codebook $\mathbb{Q}_{grid}$. By definition:
    $b_{t}=B_{ring}\quad\text{for }t\equiv 0\pmod{K}$
Consider the interval between two resets, $t\in[mK,(m+1)K]$. The bit-width initializes at $B_{ring}$ at step $mK$ and grows linearly for K1K-1 steps. The maximum width occurs at the step immediately preceding the next reset, i.e., $t_{peak}=(m+1)K-1$. Substituting into the recurrence:
$b_{peak}=B_{ring}+\sum_{i=1}^{K-1}\alpha=B_{ring}+(K-1)\alpha$
Since the reset operation strictly enforces the width back to $B_{ring}$, $b_{t}$ never exceeds $b_{peak}$.
Remark on Physical Implementation (Halo-N): In the Halo-N implementation, the "Reset Interval" $K$ is not merely a hyperparameter but is physically enforced by the hardware. The finite capacity of the registers acts as a natural boundary; when the bit-width approaches the hardware limit ($B_{max}$), it triggers an interrupt for the Ring mechanism. Thus, the finiteness of memory naturally guarantees the boundedness of the system, preventing the overflow crashes typical in standard architectures. ∎
A.2  Formal Proof of Theorem 2 (The Diophantine Error Bound)
While Theorem 1 establishes the memory boundedness of the architecture, we now verify the numerical fidelity of the Halo-N paradigm. Specifically, we prove that the Coordinate-wise Diophantine Approximation employed by the Ring mechanism yields a quantization error strictly superior to standard fixed-point arithmetic.
Theorem 2 (Precision Guarantee). Let $x\in\mathbb{R}$ be the continuous latent state of a neuron before quantization. Let $D_{max}$ be the hardware-defined denominator constraint (e.g., $2^{16}$). The Halo-N Ring mechanism guarantees that the approximation error $\epsilon$ introduced during a reset is strictly bounded by:
$\epsilon=\left|x-\frac{p}{q}\right|<\frac{1}{q\cdot D_{max}}$ (14)
Proof.
The Halo-N Ring mechanism projects the state onto the Farey Sequence via the Euclidean algorithm (conceptually equivalent to a Continued Fraction expansion). Let the continued fraction convergents of $x$ be denoted by $p_{k}/q_{k}$.
According to the property of best rational approximations (derived from Dirichlet’s Approximation Theorem), the error of the $k$-th convergent is bounded by the denominator of the next convergent:
$\left|x-\frac{p_{k}}{q_{k}}\right|<\frac{1}{q_{k}q_{k+1}}$
Our hardware allocation policy terminates the expansion at the largest index $k$ such that the denominator $q_{k}$ fits within the register limit, i.e., $q_{k}\leq D_{max}$. Consequently, the next denominator must satisfy $q_{k+1}>D_{max}$.
Substituting this inequality into the error bound yields:
$\epsilon<\frac{1}{q_{k}D_{max}}$
Implication: In the ideal case where the denominator approaches the capacity limit ($q_{k}\approx D_{max}$), the error scales as $\mathcal{O}(D_{max}^{-2})$. This represents a quadratic improvement over standard linear quantization (e.g., Fixed-Point or Integer), where the error is bounded by $\frac{1}{D_{max}}$ (i.e., $\mathcal{O}(D_{max}^{-1})$). This confirms that "Semantic Crystallization" preserves significantly higher information density per bit than floating-point truncation. ∎
@@ D.3 Silicon Feasibility: Complexity Reduction
Appendix E Appendix E: Training Dynamics and Differentiability
@@
E.1 Strategy 1: Handling Discreteness (VQ-VAE Paradigm)
@@ 2. Straight-Through Estimator (STE)
E.2 Strategy 2: Rational Mixed-Precision Training
@@
E.3 Transfer Learning: Post-Training Rationalization